OpenStudy (anonymous):

What is the exact value of the expression square root of 27 plus square root of 32 minus square root of 75? Simplify completely. 4 times the square root of 2 minus 2 times the square root of 3 square root of negative 16 16 times the square root of 2 minus 14 times the square root of 3 square root of 32 minus the square root of 12

OpenStudy (anonymous):

I don't know how much information to give you now that i got yelled at .. your answers have radical 2 and 3 in them, so try to take each one of these out of the larger radicals and simplify.

OpenStudy (anonymous):

\[\sqrt{9*3}+\sqrt{2*16}-\sqrt{25*3}\] \[3\sqrt{3}+4\sqrt{2}-5\sqrt{3}\] \[4\sqrt{2}-2\sqrt{3}\]